If I type a page in x minutes and my friend types a page in x-1 minutes how long to type 100 pages together?
Answer by jim_thompson5910(35256)

1/x + 1/(x-1) = 1/t

t(x-1) + tx = x(x-1)

t(x-1+x) = x(x-1)

t(2x-1) = x(x-1)

t = (x(x-1))/(2x-1)

So working together, you both can type out a page in (x(x-1))/(2x-1) minutes.

So 100 pages will take (100x(x-1))/(2x-1) minutes

--------------------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim
--------------------------------------------------------------------------------------------------------------